A ball is kicked from a location < 6, 0, -9> m (on the ground) with initial velocity<-13, 17.0, -4> m/s. The balls speed is low enough that air resistance is negligible (a) What is the velocity of the ball 0.6 s after being kicked? (Use the Momentum Principle!) 11.2 4 m/S (b) In this situation (constant force), which velocity will give the most accurate value for the location of the ball 0.6 s after it is kicked? the arithmetic average of the initial and final velocities the initial velocity of the ball the final velocity of the ball (c) What is the average velocity of the ball over this time interval (a vector)? -13 19.7 4 m/s (d) Use the average velocity to find the location of the ball 0.6 s after being kicked 1.8 11.82 11.4 Now consider a different time interval: the interval between the initial kick and the moment when the ball reaches its highest point. We want to find how long it takes for the ball to reach this point, and how high the ball goes (e) What is the y component of the balls velocity at the instant when the ball reaches its highest point (the end of this time interval)? m/sNow consider a different time interval: the interval between the initial kick and the moment when the ball reaches its highest point. We want to find how long it takes for the ball to reach this point, and how high the ball goes. (e) What is the y component of the balls velocity at the instant when the ball reaches its highest point (the end of this time interval)? m/s (f) Fill in the missing numbers in the equation below (update form of the Momentum Principle: mVyf-mVyì + F.iet, у//) (-mg)At (g) How long does it take for the ball to reach its highest point? (h) Knowing this time, first find the y component of the average velocity during this time interval, then use it to find the maximum height attained by the ball утахQ1. You hang a 1 kg mass from a spring and find that it stretches 0.4 m. When you hang the mass as shown from a pair of springs as shown, where each spring identical to the first one, how much does each spring stretch? A. 0.1 m B. 0.2 m C. 0.4 m D. 0.6 m E. 0.8 mQ2. You hang a 1 kg mass from a spring and find that it stretches 0.4 m. When you hang the mass as shown from a pair of springs as shown, where each spring identical to the first one, how much does the combination of springs stretch? A. 0.1 m B. 0.2 m C. 0.4 m D. 0.6 m E. 0.8 mQ3. When a mass is suspended from one spring, it stretche, the spring by distance s. Suppose it were suspended from 4 parallel strands of springs, each with 3 springs in it connected in series, what is the new extension? 4 B. s 4 D. 3s E. 4sDirection of Friction Q4. You push a block with your hand into the wall to hold it stationary. What are the direction of Fv and fFriction respectively on the block? A. Up, Left B. Up, Down C. Left, Up D. Left, Down E. None of the above Wall Block Fhandas. You push a box across the floor at a constant speed of 1 m/s, applying a horizontal constant force of magnitude 20 N. Your friend pushes the same box across the same floor at a constant speed of 2 m/s, applying a horizontal force. What is the magnitude of the force that your friend applies to the box? A. 10 N B. 20 N C. 40 N D. 80 N E. None of above
Answer #1

Solution:

Let us go to the basics of Series and parallel combinations of spring.

Q1

For a single spring

stretch = F / K = 0.4 m ....equation 1

For a parallel combination of springs, effective spring constant is given by

K' = K + K (since both springs are identical)

Thus K' = 2K

So stretch of each spring X = F/K' = F / 2K = 0.4 / 2 = 0.2 m (Answer = Option 1B)

Q2

For a single spring

stretch = F / K = 0.4 m ....equation 1

For a series combination of springs, effective spring constant is given by

1/K' = 1/K + 1/K (since both springs are identical)

Thus K' = K/2

So stretch of each spring X = F/K' = F / (K/2)

= 2F/K = 2* 0.4 = 0.8 m (Answer = Option 2E)

Thanks!!!